OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (godlovesme):

use the point slope form y-y1=m(x-x1) m is slope = -5 y1 = 4 x1= -2 plug in the info :)

OpenStudy (anonymous):

Thannkyou!

OpenStudy (godlovesme):

anytime :) thx for the medal :)